# Subdiffusive concentration for the chemical distance in Bernoulli percolation

### v1

@@ 4.1. Variance bound via entropy inequality
$\max\{\mathbb{P}({\mathrm{T}}_{z}\geq t),\mathbb{P}({\mathrm{F}}_{m}\geq t)\}\leq c^{-1}\exp\left(-\frac{ct}{\log^{2}n}\right).$ (4.41)
@@
$\mathbb{P}(|\gamma_{z}|\geq t)\leq\mathbb{P}({\mathrm{T}}_{z}\geq t)\leq c^{-1}\exp\left(-\frac{ct}{\log^{2}n}\right).$ (4.42)
@@
$\mathbb{P}(Y_{z}\geq Cn)\leq C\exp\left(-\frac{n}{C\log^{2(d+10)}n}\right),$ (4.43)
@@
$\leq Cn{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\Big{]}+({\mathbb{E}}[e^{8\lambda{{\mathrm{F}}_{m}}}])^{1/4}(\mathbb{P}(Y_{z}\geq Cn))^{1/4}({\mathbb{E}}[Y_{z}^{2}])^{1/2},$ (4.44)
thanks to Cauchy-Schwarz inequality. By (4.42) and Corollary 3.4 (i),
${\mathbb{E}}[Y_{z}^{2}]\leq{\mathcal{O}}(n^{2}).$ (4.45)
It follows from (4.41) that for all $\lambda\leq\frac{1}{\log^{2(d+11)}n}$,
@@
where $C^{\prime}$ is the constant in (4.41). Putting this estimate, (4.43), (4.4), and (4.45) together, we have for all $\lambda\leq\frac{1}{\log^{2(d+11)}n}$,
${\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}Y_{z}{\mathbb{I}}(Y_{z}\geq Cn)\Big{]}\leq C\exp\left(-\frac{n}{2C\log^{2(d+10)}n}\right).$ (4.46)
Thanks to (4.41) again, we have for all $-\frac{1}{\log^{2(d+11)}n}\leq\lambda<0$,
@@
with $C^{\prime}$ being the constant defined as in (4.41). Combining the last two display equations with (4.4) yields for all $|\lambda|\leq\frac{1}{\log^{2(d+11)}n}$,
@@
which together with (4.40) implies the desired result. ∎
5. Comparison of the graph distance and the first truncated passage time
In order to prove Theorem 1.4, our aim is to upper bound the difference between ${\mathrm{D}}^{*}_{n}$ and ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})$ by the total size of a collection of bypasses avoiding all closed ($\log^{2}n$-weight) edges along $\gamma$-a geodesic of ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})$. More precisely, we introduce an inductive process that reveals a set of closed edges $\Gamma\subset{\mathrm{clo}}(\gamma)$ that are separable, as described by property (ii) below. Consequently, the discrepancy can be bounded via the total sum of effective radii associated with edges in $\Gamma$. Recall that $C_{*}$ be the constant and $(R_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ be the effective radii defined as in Section 2.
@@ 5. Comparison of the graph distance and the first truncated passage time
Let $\gamma$ be a geodesic of ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})$. Then on the event $\{\{0^{*},(n{\mathbf{e}}_{1})^{*}\}\not\subset\Lambda_{3R_{e}}(e)\,\,\forall e\in\mathrm{clo}(\gamma)\}$, there exists a subset $\Gamma\subset{\mathrm{clo}}(\gamma)$ such that
@@
Let us first show that the following procedure (G) is well defined:
Input: A path $\eta\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$ satisfying ${\mathrm{clo}}(\eta)\neq\varnothing$ and $\{0^{*},(n{\mathbf{e}}_{1})^{*}\}\not\subset\Lambda_{3R_{e}}(e)\,\,\forall e\in\mathrm{clo}(\eta)\}$.
@@
    $\text{clo}(\eta_{e_{*}})\cap\Lambda_{R_{e_{*}}}({e_{*}})=\varnothing$;
@@
Indeed, we simply take $e_{*}\in{\mathrm{clo}}(\gamma)$ with the largest effective radius, using a rule to break ties when there are several maximizing edges. Applying Proposition 2.6 to $\eta\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$ and $e_{*}$, we obtain $\eta_{e_{*}}\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$, a modified path of $\eta$, satisfying $({\rm c})$, $({\rm d})$ and that $\eta_{e_{*}}\setminus\eta$ is open, and $|\eta_{e_{*}}\setminus\eta|\leq 2C_{*}R_{e_{*}}$. Therefore,
${\mathrm{T}}(\eta)\leq{\mathrm{T}}(\eta_{e_{*}})\leq{\mathrm{T}}(\eta)+{\mathrm{T}}(\eta_{e_{*}}\setminus\eta)-t_{e_{*}}\leq{\mathrm{T}}(\eta)+2C_{*}R_{e_{*}}-\log^{2}n,$
which implies $R_{e_{*}}\geq\log^{2}n/2C_{*}$ and thus (b) follows.
Figure 4. Illustration for covering process of all closed edges in the original geodesic $\gamma$
We set $\eta^{0}=\gamma$ and iteratively apply the procedure (G) to obtain a sequence of edges $(e_{k})_{k=1}^{\ell}\subset{\mathrm{clo}}(\eta^{0})$ and paths $(\eta^{k})_{k=1}^{\ell}$ such that $e_{k}=e_{*}(\eta^{k-1})$, $\eta^{k}=(\eta^{k})_{e_{k}}\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$ for $k=1,\ldots,\ell$, and that (a)-(d) hold for each pair of $(e_{k},\eta^{k})_{k=1}^{\ell}$ and more importantly $\eta^{\ell}$ is open (see Figure 4 for illustration). Remark that the iteration stops after some finite steps $\ell$, since $|{\mathrm{clo}}(\eta^{0})|=|{\mathrm{clo}}(\gamma)|$ is finite and $|{\mathrm{clo}}(\eta^{k})|<|{\mathrm{clo}}(\eta^{k-1})|$ using (b) and (c). Setting $\Gamma=\{e_{1},\ldots,e_{\ell}\}$, we can easily deduce (i)–(iii) from (a)–(d). Indeed, (i) follows from (b). By (a) and (b), the sequence $(R_{e_{k}})_{k=1}^{\ell}$ is decreasing and hence by (c)
$\forall 1\leq i<j\leq\ell,\quad\|e_{i}-e_{j}\|_{\infty}\geq R_{e_{i}}=\max\{R_{e_{i}},R_{e_{j}}\}.$
Since $\eta^{\ell}$ is an open path from $0^{*}$ to $(ne_{1})^{*}$, the property (${\rm d}$) shows
@@
$\mathbb{P}\left(|{{\mathrm{D}}^{*}_{n}}-{\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})|\geq L\right)\leq c^{-1}\exp(-cL/\log L).$ (5.1)
Indeed, by (5.1) and Lemma 2.8,
@@
where $c_{1}$ is a positive constant. The remainder of this section is devoted to proving (5.1). Let $\gamma_{n}$ be a geodesic of ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})$. By Lemma 2.8 and Lemma 2.10, there exist a positive constant $C$ such that
$\mathbb{P}({\mathrm{D}}^{*}_{n}\leq Cn/2;\|0-0^{*}\|_{\infty},\|n{\mathbf{e}}_{1}-(n{\mathbf{e}}_{1})^{*}\|_{\infty}\leq n/4)\geq 1-C\exp(-n/C).$ (5.2)
Remark further that ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})\leq{\mathrm{D}}^{*}_{n}$, and if ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})\leq k$ then $\gamma_{n}\subset\Lambda_{k}(0^{*})$, since $t_{e}\geq 1$ for all $e$. Hence, if the event in (5.2) holds then $\gamma_{n}\subset\Lambda_{Cn}$, and thus we have
$\mathbb{P}({\mathcal{E}})\geq 1-C\exp(-n/C),\qquad{\mathcal{E}}:=\{\gamma_{n}\subset\Lambda_{Cn}\}.$ (5.3)
@@
:=${\mathcal{E}}_{*}:=\{\{0^{*},(n{\mathbf{e}}_{1})^{*}\}\not\subset\Lambda_{3R_{e}}(e)\,\forall\,e\in\gamma_{n}\}.$
Using (5.2), (5.3), and Proposition 2.4, there are positive constants $C$ and $c$ such that
@@
+($\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\mathbb{P}(\{\|0-0^{*}\|_{\infty}+\|n{\mathbf{e}}_{1}-(n{\mathbf{e}}_{1})^{*}\|_{\infty}\geq n/2)$
$\leq{\mathbb{P}}(\exists\,e\in\Lambda_{Cn}:R_{e}\geq n/12)+3C\exp(-cn)\leq 4C\exp(-cn).$ (5.4)
@@
Case 2: $L<Cn$. Using (5.3), (5), Propositions 2.4 and 5.1,
@@
[…]
@@
    $\|e-e^{\prime}\|_{\infty}\geq\max\{R_{e},R_{e^{\prime}}\}$, for all $e,e^{\prime}\in\Gamma_{n}$,
@@
Therefore, if (c) occurs then there exists $0\leq q\leq a_{n}$ such that $N_{q}\geq B_{q}$ where for $1\leq q\leq a_{n}$,
@@
$\mathbb{P}\left(\exists\,\Gamma_{n}\subset\Lambda_{Cn}\textrm{ satisfying (a)--(c)}\right)\leq\sum_{q=0}^{a_{n}}\mathbb{P}\left(\exists\,\Gamma_{n}\subset\Lambda_{Cn}\textrm{ satisfying (a), (b)};\,N_{q}\geq B_{q}\right).$ (5.6)
Moreover, by (b), if $R_{e}$ and $R_{e^{\prime}}$ are in $[M_{q},2M_{q}]$ for some $e,e^{\prime}\in\Gamma_{n}$, then
$\|e-e^{\prime}\|_{\infty}\geq\max\{R_{e},R_{e^{\prime}}\}\geq M_{q}.$ (5.7)
@@
$\mathbb{P}\left(\exists\,\Gamma_{n}\subset\Lambda_{Cn}\textrm{ satisfying (a), (b)};\,N_{q}\geq B_{q}\right)\leq S_{q},$ (5.8)
@@
$S_{q}:=$ ($\mathbb{P}\Big{(}\,\exists\,\{e_{1},\ldots,e_{B_{q}}\}\subset\Lambda_{Cn}:R_{e_{j}}\in[M_{q},2M_{q}]\,\forall\,j\in[B_{q}];\|e_{j}-e_{k}\|_{\infty}\geq M_{q}\,\forall j\neq k\in[B_{q}]\Big{)},$ (5.9)
here recall the notation $[a]=\{1,\ldots,a\}$. The following claim is straightforward.
@@
$\|e-e^{\prime}\|_{\infty}\geq M,\quad\forall e,e^{\prime}\in E,$
we can find $E^{\prime}\subset E$ such that $|E^{\prime}|\geq c|E|$ and $\|e-e^{\prime}\|_{\infty}\geq 5CM$ for all $e,e^{\prime}\in E^{\prime}$.
By this claim, there exists a positive constant $c$ such that for any $1\leq q\leq a_{n}$ if the event in (5.9) occurs then we can find $E^{\prime}\subset\{e_{1},\ldots,e_{B_{q}}\}$ such that $|E^{\prime}|\geq\lfloor cB_{q}\rfloor$ and $\|e-e^{\prime}\|_{\infty}\geq 5C_{*}M_{q}$ for all $e,e^{\prime}\in E^{\prime}$. As a result, we have
$S_{q}$ ($\leq\mathbb{P}\left(\,\exists\,\{e^{\prime}_{1},\ldots,e^{\prime}_{C_{q}}\}\in\mathcal{T}_{q}:R_{e^{\prime}_{j}}\in[M_{q},2M_{q}]\,\forall j\in[C_{q}]\right)$
$\leq\sum_{\{e^{\prime}_{1},\ldots,e^{\prime}_{C_{q}}\}\in\mathcal{T}_{q}}\mathbb{P}\left(R_{e^{\prime}_{j}}\in[M_{q},2M_{q}]\,\forall j\in[C_{q}]\right),$ (5.10)
@@
We remark that the event $R_{e}\in[M_{q},2M_{q}]$ only depends on the state of edges in $\Lambda_{2C_{*}M_{q}}(e)$. Therefore, given $\{e^{\prime}_{1},\ldots,e^{\prime}_{C_{q}}\}\in\mathcal{T}_{q}$, the family of the events ($(\{R_{e^{\prime}_{j}}\in[M_{q},2M_{q}]\})_{i\in[C_{q}]}$ are independent. Hence,
@@
where $\alpha$ is a positive constant. Here for the inequality we have used Proposition 2.4 with the remark that $M_{q}\leq M_{a_{n}}\leq n^{2}$. This estimate together with (5) yields for all $1\leq q\leq a_{n}$,
@@
Combining the above estimate with (5.8), (5.6), and (5), we obtain (5.1). ∎
Acknowledgment. We would like to thank Shuta Nakajima for his valuable comments. V. Q. Nguyen is partially supported by the Master, PhD Scholarship Programme of Vingroup Innovation Foundation (VINIF).

### v2

@@ 4.1. Variance bound via entropy inequality
$\max\{\mathbb{P}({\mathrm{T}}^{z}\geq t),\mathbb{P}({\mathrm{F}}_{m}\geq t)\}\leq c^{-1}\exp\left(-\frac{ct}{\log^{2}n}\right).$ (82)
@@
$\mathbb{P}(|\gamma_{z}|\geq t)\leq\mathbb{P}({\mathrm{T}}^{z}\geq t)\leq c^{-1}\exp\left(-\frac{ct}{\log^{2}n}\right).$ (83)
@@
$\mathbb{P}(Y_{z}\geq Cn)\leq C\exp\left(-\frac{n}{C\log^{2(d+10)}n}\right),$ (84)
@@
$\leq Cn{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\Big{]}+({\mathbb{E}}[e^{8\lambda{{\mathrm{F}}_{m}}}])^{1/4}(\mathbb{P}(Y_{z}\geq Cn))^{1/4}({\mathbb{E}}[Y_{z}^{2}])^{1/2},$ (85)
thanks to Cauchy-Schwarz inequality. By (83) and Corollary 3.4 (i),
${\mathbb{E}}[Y_{z}^{2}]\leq{\mathcal{O}}(n^{2}).$ (86)
It follows from (82) that for all $\lambda\leq\frac{1}{\log^{2(d+11)}n}$,
@@
where $C^{\prime}$ is the constant in (82). Putting this estimate, (84), (4.4), and (86) together, we have for all $\lambda\leq\frac{1}{\log^{2(d+11)}n}$,
${\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}Y_{z}{\mathbb{I}}(Y_{z}\geq Cn)\Big{]}\leq C\exp\left(-\frac{n}{2C\log^{2(d+10)}n}\right).$ (87)
Thanks to (82) again, we have for all $\frac{1}{\log^{2(d+11)}n}\leq\lambda<0$,
@@
with $C^{\prime}$ being the constant defined as in (82). Combining the last two display equations with (4.4) yields for all $|\lambda|\leq\frac{1}{\log^{2(d+11)}n}$,
@@
which together with (81) implies the desired result. ∎
5. Comparison of the graph distance and the truncated first passage time
In order to prove Theorem 1.4, our aim is to upper bound the difference between ${\mathrm{D}}^{*}_{n}$ and ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})$ by the total size of a collection of bypasses avoiding all closed ($\log^{2}n$-weight) edges along $\gamma$–a geodesic of ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})$. More precisely, we introduce an inductive process that reveals a set of closed edges $\Gamma\subset{\mathrm{clo}}(\gamma)$ that are separable, as described by property $\rm(ii)$} below. Consequently, the discrepancy can be bounded via the total sum of effective radii associated with edges in $\Gamma$. Recall that $C_{*}$ be the constant and $(R_{e})_{e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}}$ be the effective radii defined as in Section 2.
@@ 5. Comparison of the graph distance and the truncated first passage time
Let $\gamma$ be a geodesic of ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})$. Suppose that $\{0^{*},(n{\mathbf{e}}_{1})^{*}\}\not\subset\Lambda_{3R_{e}}(e)$ for all $e\in\mathrm{clo}(\gamma)$. Then the following holds.
@@
    There exists a subset $\Gamma\subset{\mathrm{clo}}(\gamma)$ such that
    1. (iia)
@@
If $e\in{\mathrm{clo}}(\gamma)$ satisfies $\{0^{*},(n{\mathbf{e}}_{1})^{*}\}\not\subset\Lambda_{3R_{e}}(e)$, we can apply Proposition 2.8 to $\gamma\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$ and $e$, obtaining another path $\gamma_{e}\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$ such that $\gamma_{e}\setminus\gamma$ is open, and $|\gamma_{e}\setminus\gamma|\leq 2C_{*}R_{e}$. Hence,
${\mathrm{T}}(\gamma)\leq{\mathrm{T}}(\gamma_{e})\leq{\mathrm{T}}(\gamma)+{\mathrm{T}}(\gamma_{e}\setminus\gamma)-t_{e}\leq{\mathrm{T}}(\gamma)+2C_{*}R_{e}-\log^{2}n,$
which implies that $R_{e}\geq\log^{2}n/2C_{*}$ and hence $\rm(i)$} follows.
It remains to prove $\rm(ii)$}. Let us first show that the following procedure (G) is well defined:
Input: A path $\eta\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$ satisfying ${\mathrm{clo}}(\eta)\neq\varnothing$ and $\{0^{*},(n{\mathbf{e}}_{1})^{*}\}\not\subset\Lambda_{3R_{e}}(e)\,\,\forall e\in\mathrm{clo}(\eta)$.
@@
    ${\mathrm{clo}}(\eta_{e_{*}})\cap\Lambda_{R_{e_{*}}}({e_{*}})=\varnothing$;
@@
Indeed, we simply take $e_{*}\in{\mathrm{clo}}(\gamma)$ with the largest effective radius, using a rule to break ties when there are several maximizing edges. Applying Proposition 2.8 to $\eta\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$ and $e_{*}$, we obtain $\eta_{e_{*}}\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$, a modified path of $\eta$, satisfying $({\rm b}),({\rm c})$, $({\rm d})$.
Figure 5. Illustration for covering process of all closed edges in the original geodesic $\gamma$. Black lines (resp. red) represent closed edges (resp. open) in $\gamma$, and the open paths $(\eta^{k}\setminus\eta^{k-1})_{k=1}^{\ell}$ are shown by blue.
We set $\eta^{0}=\gamma$ and iteratively apply the procedure (G) to obtain a sequence of edges $(e_{k})_{k=1}^{\ell}\subset{\mathrm{clo}}(\eta^{0})$ and paths $(\eta^{k})_{k=1}^{\ell}$ such that $e_{k}=e_{*}(\eta^{k-1})$, $\eta^{k}=(\eta^{k})_{e_{k}}\in\mathbb{G}_{*}(0^{*},(n{\mathbf{e}}_{1})^{*})$ for $k=1,\ldots,\ell$, and that $\rm(a)$$\rm(d)$} hold for each pair of $(e_{k},\eta^{k})_{k=1}^{\ell}$ and more importantly $\eta^{\ell}$ is open (see Figure 5 for illustration). Remark that the iteration stops after some finite steps $\ell$, since $|{\mathrm{clo}}(\eta^{0})|=|{\mathrm{clo}}(\gamma)|$ is finite and $|{\mathrm{clo}}(\eta^{k})|<|{\mathrm{clo}}(\eta^{k-1})|$ using (b) and (c). Setting $\Gamma=\{e_{1},\ldots,e_{\ell}\}$, we can easily deduce $\rm(iia)$} and $\rm(iib)$} from $\rm(a)$$\rm(d)$}. Indeed, by $\rm(a)$} and $\rm(b)$}, the sequence $(R_{e_{k}})_{k=1}^{\ell}$ is decreasing and hence by $\rm(c)$},
$\forall 1\leq i<j\leq\ell,\quad\|e_{i}-e_{j}\|_{\infty}\geq R_{e_{i}}=\max\{R_{e_{i}},R_{e_{j}}\},$
which follows $\rm(iia)$}. Since $\eta^{\ell}$ is an open path from $0^{*}$ to $(ne_{1})^{*}$, the property ($\rm d$}) shows
@@
$\mathbb{P}\left(|{{\mathrm{D}}^{*}_{n}}-{\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})|\geq L\right)\leq c^{-1}\exp(-cL/\log L).$ (88)
Indeed, by (88) and Lemma 2.10,
@@
where $c_{1}$ is a positive constant. The remainder of this section is devoted to proving (88). Let $\gamma_{n}$ be a geodesic of ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})$ with some deterministic rules breaking ties. By Lemma 2.10 and Lemma 2.12, there exist a positive constant $C$ such that
$\mathbb{P}({\mathrm{D}}^{*}_{n}\leq Cn/2;\|0-0^{*}\|_{\infty},\|n{\mathbf{e}}_{1}-(n{\mathbf{e}}_{1})^{*}\|_{\infty}\leq n/4)\geq 1-C\exp(-n/C).$ (89)
Remark further that ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})\leq{\mathrm{D}}^{*}_{n}$, and if ${\mathrm{T}}(0^{*},(n{\mathbf{e}}_{1})^{*})\leq k$ then $\gamma_{n}\subset\Lambda_{k}(0^{*})$, since $t_{e}\geq 1$ for all $e$. Hence, if the event in (89) holds then $\gamma_{n}\subset\Lambda_{Cn}$, and thus we have
$\mathbb{P}({\mathcal{E}})\geq 1-C\exp(-n/C),\qquad{\mathcal{E}}:=\{\gamma_{n}\subset\Lambda_{Cn}\}.$ (90)
@@
${\mathcal{E}}_{*}:=\{\{0^{*},(n{\mathbf{e}}_{1})^{*}\}\not\subset\Lambda_{3R_{e}}(e)\,\forall\,e\in\gamma_{n}\}.$
Using (89), (90), and Proposition 2.6, there are positive constants $C$ and $c$ such that
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\mathbb{P}(\{\|0-0^{*}\|_{\infty}+\|n{\mathbf{e}}_{1}-(n{\mathbf{e}}_{1})^{*}\|_{\infty}\geq n/2)$
$\leq{\mathbb{P}}(\exists\,e\in\Lambda_{Cn}:R_{e}\geq n/12)+3C\exp(-cn)\leq 4C\exp(-cn).$ (91)
@@
Case 2: $L<Cn$. Using (90), (5), Propositions 2.6 and 5.1,
@@
$\leq\mathbb{P}({\mathcal{E}}^{c})+\mathbb{P}({\mathcal{E}}_{*}^{c})+\mathbb{P}\left(\exists\,e\in\Lambda_{Cn}:R_{e}\geq L\right)+\mathbb{P}\left(\exists\,\Gamma_{n}\subset\Lambda_{Cn}\textrm{ satisfying (a)--(c)}\right)$
@@
    $\|e-e^{\prime}\|_{\infty}\geq\max\{R_{e},R_{e^{\prime}}\}$, for all $e\neq e^{\prime}\in\Gamma_{n}$,
@@
Therefore, if (c) occurs then there exists $0\leq q\leq a_{n}$ such that $N_{q}\geq B_{q}$ where for each $0\leq q\leq a_{n}$,
@@
$\mathbb{P}\left(\exists\,\Gamma_{n}\subset\Lambda_{Cn}\textrm{ satisfying (a)--(c)}\right)\leq\sum_{q=0}^{a_{n}}\mathbb{P}\left(\exists\,\Gamma_{n}\subset\Lambda_{Cn}\textrm{ satisfying (a), (b)};\,N_{q}\geq B_{q}\right).$ (93)
Moreover, by (b), if $R_{e}$ and $R_{e^{\prime}}$ are in $M_{q},2M_{q}]$ for some $e\neq e^{\prime}\in\Gamma_{n}$, then
$\|e-e^{\prime}\|_{\infty}\geq\max\{R_{e},R_{e^{\prime}}\}\geq M_{q}.$ (94)
@@
$\mathbb{P}\left(\exists\,\Gamma_{n}\subset\Lambda_{Cn}\textrm{ satisfying (a), (b)};\,N_{q}\geq B_{q}\right)\leq S_{q},$ (95)
@@
$S_{q}:=$ $\mathbb{P}\Big{(}\,\exists\,\{e_{1},\ldots,e_{B_{q}}\}\subset\Lambda_{Cn}:R_{e_{j}}\in[M_{q},2M_{q}]\,\forall\,j\in[B_{q}];\|e_{j}-e_{k}\|_{\infty}\geq M_{q}\,\forall j\neq k\in[B_{q}]\Big{)},$ (96)
here recall the notation $a]=\{1,\ldots,a\}$. The following claim is straightforward.
@@
$\|e-e^{\prime}\|_{\infty}\geq M,\quad\forall e\neq e^{\prime}\in E,$
we can find $E^{\prime}\subset E$ such that $|E^{\prime}|\geq c|E|$ and $\|e-e^{\prime}\|_{\infty}\geq 5CM$ for all $e\neq e^{\prime}\in E^{\prime}$.
By this claim, there exists a positive constant $c$ such that for any $0\leq q\leq a_{n}$ if the event in (96) occurs then we can find $E^{\prime}\subset\{e_{1},\ldots,e_{B_{q}}\}$ such that $|E^{\prime}|\geq\lfloor cB_{q}\rfloor$ and $\|e-e^{\prime}\|_{\infty}\geq 5C_{*}M_{q}$ for all $e\neq e^{\prime}\in E^{\prime}$. As a result, we have
$S_{q}$ $\leq\mathbb{P}\left(\,\exists\,\{e^{\prime}_{1},\ldots,e^{\prime}_{C_{q}}\}\in\mathcal{T}_{q}:R_{e^{\prime}_{j}}\in[M_{q},2M_{q}]\,\forall j\in[C_{q}]\right)$
$\leq\sum_{\{e^{\prime}_{1},\ldots,e^{\prime}_{C_{q}}\}\in\mathcal{T}_{q}}\mathbb{P}\left(R_{e^{\prime}_{j}}\in[M_{q},2M_{q}]\,\forall j\in[C_{q}]\right),$ (97)
@@
We remark that the event $R_{e}\in[M_{q},2M_{q}]$ only depends on the state of edges in $\Lambda_{2C_{*}M_{q}}(e)$. Therefore, given $\{e^{\prime}_{1},\ldots,e^{\prime}_{C_{q}}\}\in\mathcal{T}_{q}$, the family of the events $(\{R_{e^{\prime}_{j}}\in[M_{q},2M_{q}]\})_{i\in[C_{q}]}$ are independent. Hence,
@@
where $\alpha$ is a positive constant. Here for the inequality we have used Proposition 2.6 with the remark that $M_{q}\leq M_{a_{n}}\leq n^{2}$. This estimate together with (5) yields for all $0\leq q\leq a_{n}$,
@@
Combining the above estimate with (95), (93), and (5), we obtain (88). ∎
Acknowledgment. We would like to thank Shuta Nakajima for his valuable comments. V. Q. Nguyen was funded by the Master, PhD Scholarship Programme of Vingroup Innovation Foundation (VINIF), code VINIF.2024.TS.002.